OpenStudy (anonymous):

Write the following as an equation, using x for the unknown number. Then solve. The sum of three consecutive integers is 483. Find the numbers.

OpenStudy (anonymous):

Let the three integers be (x-1), x and (x+1)

OpenStudy (anonymous):

answer is 161

OpenStudy (anonymous):

their sum will be x-1 + x + x+1 = 483

OpenStudy (anonymous):

this will give you 3x = 483 and x = 161

OpenStudy (anonymous):

so the three integers will be 160, 161 and 162

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

how did you get 161? @akitav

OpenStudy (anonymous):

we got 3x= 483. divide both sides by 3 you'll get \[\frac{3x}{3}=\frac{483}{3}\]

OpenStudy (anonymous):

the LHS becomes x and the RHS becomes 161 (483 divided by 3) so we get x = 161

OpenStudy (anonymous):

thank you! @akitav

OpenStudy (anonymous):

Its been a pleasure :)